# Real-World Adversarial Attacks on RF-Based Drone Detectors

### v3

@@ 1 Introduction
We introduce the first physical adversarial attack targeting drone transmission specifically designed for RF based OD models. Rather than optimizing perturbations in the digital feature space, our approach generates adversarial waveforms directly in the I/Q domain and emits them over-the-air (OTA) from an independent transmitter. This creates interference that is combined with legitimate drone signals but eliminates the need for the complex digital-to-RF conversion and ensures compatibility with standard RF hardware. We perform both digital and physical evaluation using the same optimized I/Q-domain CUAP across various OD architectures across various OD architectures, demonstrating the generalizability of I/Q domain adversarial perturbation in RF detection systems.
@@ 3.2 Optimization Objective
Suppressing the detection of the target class. To selectively suppress the target class $c^{*}$ and avoid attacking false positives, we retain only detections that spatially align with the target class ground truth: Let $\mathcal{I}=\{1,\dots,N\}$ index all pre-NMS predictions of the detector, where each prediction $(b_{i},s_{i,1},\ldots,s_{i,C})$ contains the bounding box $b_{i}$ and the confidence score $s_{i,c}$ for each class $c\in C$, and let $\mathcal{G}_{c^{*}}=\{g_{k}\}_{k=1}^{K_{c^{*}}}\!$ be the ground-truth boxes of the target class $c^{*}$. We define the set of predictions that match the ground truth using Intersection-over-union (IoU) threshold $\tau=0.5$:
@@
$\mathcal{L}_{\text{evade}}=\sum_{k}\sum_{i\in\mathcal{A}_{k}}-\log\!\bigl(1-s_{i,c^{*}}\bigr)$ (3)
@@
$\mathcal{L}_{\text{protect}}=\frac{1}{NC^{\prime}}\sum_{i=1}^{N}\sum_{\begin{subarray}{c}c=1\\ c\neq c^{*}\end{subarray}}^{C}\big\|s_{i,c}(x_{\text{adv}})-s_{i,c}(x_{\text{clean}})\big\|$ (4)
@@ 4.1 Dataset
Attack set generation. We assume an adversary that aims to prevent their drone’s detection by OD systems. They can control their drone’s transmit channel (i.e., center frequency) but have no control of other drones’ transmission channels. Accordingly, we recorded the target drone OTA in a realistic, noisy environment; other drones’ signals were digitally mixed into the I/Q stream at random channels across the 2.4 GHz ISM band. This setup reflects a realistic threat model in which the adversary can manipulate the target drone’s emissions but cannot influence coexisting signals. The attack set used to train the adversarial waveform was split into 70%, 15%, and 15% for the training, validation, and test subsets, respectively.
@@ 4.4 Experimental Results
Digital attacks. For the digital evaluation, we digitally injected the CUAP into the I/Q stream prior to spectrogram formation (Eq. 1). As baselines we report the results obtained on clean I/Q signals and with random AWGN. We considered the following four attacker-knowledge scenarios: White-box (WB): per-architecture CUAP with full access to the target detector’s weights; Gray-box (GB): CUAP trained on a surrogate with the same architecture using a disjoint training subset; Closed-set (CS): CUAP trained jointly on all surrogates; Black-box (BB): leave-one-architecture-out training where the held-out architecture serves as the BB target.

### v4

@@ 1 Introduction
We introduce the first physical adversarial attack targeting drone transmission specifically designed for RF based OD models. Rather than optimizing perturbations in the digital feature space, our approach generates adversarial waveforms directly in the I/Q domain and emits them over-the-air (OTA) from an independent transmitter. This creates interference that is combined with legitimate drone signals but eliminates the need for the complex digital-to-RF conversion and ensures compatibility with standard RF hardware. We perform both digital and physical evaluation using the same optimized I/Q-domain CUAP across various OD architectures, demonstrating the generalizability of I/Q domain adversarial perturbation in RF detection systems.
@@ 3.2 Optimization Objective
Suppressing the detection of the target class. To selectively suppress the target class $c^{*}$ and avoid attacking false positives, we retain only detections that spatially align with the target class ground truth: Let $\mathcal{I}=\{1,\dots,N\}$ index all pre-NMS predictions of the detector, where each prediction $(b_{i},s_{i,1},\ldots,s_{i,C})$ contains the bounding box $b_{i}$ and the confidence score $s_{i,c}$ for each class $c\in C$, and let $\mathcal{G}_{c^{*}}$ be the ground-truth boxes of the target class $c^{*}$. We define the set of predictions that match the ground truth using intersection-over-union (IoU) threshold $\tau=0.5$:
@@
$\mathcal{L}_{\text{evade}}=\frac{1}{|\mathcal{S}_{c^{*}}|}\sum_{i\in\mathcal{S}_{c^{*}}}-\log\bigl(1-s_{i,c^{*}}\bigr)$ (3)
@@
$\mathcal{L}_{\text{protect}}=\frac{1}{N(C-1)}\sum_{i=1}^{N}\sum_{\begin{subarray}{c}c=1\\ c\neq c^{*}\end{subarray}}^{C}\big\|s_{i,c}(x_{\text{adv}})-s_{i,c}(x_{\text{clean}})\big\|$ (4)
@@ 4.1 Dataset
Attack set generation. We assume an adversary that aims to prevent their drone’s detection by OD systems. They can control their drone’s transmit channel (i.e., center frequency) but have no control of other drones’ transmission channels. Accordingly, we recorded the target drone OTA in a realistic, noisy environment; other drones’ signals were digitally mixed into the I/Q stream at random channels across the 2.4 GHz ISM band. This setup reflects a realistic threat model in which the adversary transmits an additional OTA adversarial signal but cannot influence coexisting signals. The attack set used to train the adversarial waveform was split into 70%, 15%, and 15% for the training, validation, and test subsets, respectively.
@@ 4.4 Experimental Results
Digital attacks. For the digital evaluation, we digitally injected the CUAP into the I/Q stream prior to spectrogram formation (Eq. 1). As baselines we report the results obtained on clean I/Q signals and with random AWGN. We considered the following four attacker-knowledge scenarios: White-box (WB): per-architecture CUAP with full access to the target detector’s weights; Gray-box (GB): CUAP optimized on a surrogate with the same architecture using a disjoint training subset; Closed-set (CS): CUAP optimized jointly on all surrogates; Black-box (BB): leave-one-architecture-out training where the held-out architecture serves as the BB target.